40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 11 23:31:50 2017, johannes, Update, Cameras, post-vent camera capture comparison 13x
    Reply  Tue Sep 26 22:11:08 2017, johannes, Update, Cameras, post-vent camera capture comparison lens_distance.pdfImage__2017-09-26__18-54-29.jpgIMG_20170926_190157222_HDR.jpgIMG_20170926_185649325.jpg
       Reply  Mon Oct 2 12:44:45 2017, johannes, Update, Cameras, Basler 120gm calibration calib_20170930_152.pdfcalib_20170930_153.pdf
          Reply  Mon Oct 2 23:16:05 2017, gautam, HowTo, Cameras, CCD calibration 
             Reply  Tue Oct 3 01:58:32 2017, johannes, HowTo, Cameras, CCD calibration gige_calibration.pdf
                Reply  Thu Oct 12 01:03:49 2017, johannes, HowTo, Cameras, ETMX GigE side view IMG_20171011_164549698.jpgImage__2017-10-11__16-52-01.pngGigE_lens_position_helper.py.zip
                   Reply  Thu Oct 12 07:56:33 2017, Steve, HowTo, Cameras, ETMX GigE side view at 50 deg of IR scattering Image__2017-10-11__15-29-52_15k400g.pngImage__2017-10-12__15-50-18wipedRefocud2.pngcamCan16cm.jpg
                      Reply  Wed Oct 18 11:37:58 2017, johannes, HowTo, Cameras, ETMX GigE side view at 50 deg 
                Reply  Wed Oct 18 15:26:58 2017, johannes, HowTo, Cameras, Revision: CCD calibration calib_20170930_152.pdfcalib_20170930_153.pdf
Message ID: 13348     Entry time: Mon Oct 2 12:44:45 2017     In reply to: 13334     Reply to this: 13352
Author: johannes 
Type: Update 
Category: Cameras 
Subject: Basler 120gm calibration 

Disclaimer: Wrong calibration factors! See https://nodus.ligo.caltech.edu:8081/40m/13391

The two acA640-120gm Basler GigE-Cams have been calibrated. I used the collimated output of a fiber that carried the auxiliary laser light from the PSL table. With a non-polarizing beam splitter some of the light was picked off onto a PD, and I modified the RF amplitude of the AOM drive signal to vary the power coming out of the fiber. The fiber output was directed at a white paper, which was placed 1.06m from the front of the lens tube assembly, which is where the focal plane is. Using the Pylon Viewer App I made sure that the entirety of the beam spot was imaged onto the CCD. Since the camera sensor is 1/4" across, I removed the camera from the lens tube and instead placed the Ophir power meter head at the position of the sensor and measured the power reported versus PD voltage, which turned out to be 1.5 V/uW.

The camera was put back in place and I used the Pypylon package Gautam had stumbled upon to sweep the exposure time from 100us to 10ms at different light power settings including no laser light at all for background subtraction, and rather than keeping the full bitmap data for O(100s) of images I recorded only the quantities

  1. Pixel Max
  2. Pixel Sum
  3. Pixel Mean
  4. Pixel Standard Deviation
  5. Pixel Median

I performed this procedure for both the 152 and 153 cameras and plotted the pixel sum and the pixel max vs the exposure time. All the exposures were taken at a gain setting of 100, which is the smallest possible setting (out of 100-600). To obtain the calibration factor I use the input power Pin=75nW in the 'safe' region 1ms to 10ms where the pixel sum looks smooth and the CCD is reportedly not saturated.

Camera IP Calibration Factor CF
192.168.113.152 8.58 W*s
192.168.113.153 7.83 W*s

The incident power can be calculated as Pin =CF*Total(Counts-DarkCounts)/ExposureTime.

Attachment 1: calib_20170930_152.pdf  24 kB  | Show | Hide all | Show all
Attachment 2: calib_20170930_153.pdf  24 kB  | Hide | Hide all | Show all
calib_20170930_153.pdf
ELOG V3.1.3-